# Evaluating crop rotations around the world using satellite imagery and causal machine learning

### v1

@@ 3.2 Discussion of heterogeneity with weather findings
While these findings of how rotation effects vary with weather are consistent with the existing literature (Appendix J), we note that our findings go far beyond the scope of most prior work. In particular, existing claims in the experimental literature looked at weather interactions after aggregating over many different precrop and outcome crop combinations, due to limitations in the number and types of experiments [10, 5, 9, 8]. Such claims can therefore reflect how precrops and outcome crops in experiments vary across regions with differing climates rather than how the benefit of rotation varies with weather. In addition, claims in the observational literature about how rotation benefits varied with whether all focused on a single country and only one of them provided uncertainty quantifications for these claims [12, 17, 19, 20] (Table A5).
@@ 3.3 Discussion of diversification benefit findings
Our results suggest that diversification beyond a simple rotation has diminishing returns (see Table A8 or Figures 3 and 5). In particular, we find that the average benefit of a 2-crop rotation across our 36 rotation-country pairs is about 3.7 times larger than the average additional benefit of switching from 2-crop rotations to a more diverse sequence. This finding agrees with a recent metanalysis of 45 experiments in China [9] that found crop rotation increased yields by 20 when compared to monoculture on average but extended (diversified) rotations provided only an additional 4 yield boost. Similarly, a recent global metanalysis of 462 crop rotation experiments involving legumes found that the inclusion of an additional legume precrop in a rotation had lower benefits when the control rotation was already diverse [5]. A recent analysis of field experiments in the European Union and Africa [39] found no additional benefits from increased diversification beyond a switch from monocropping to simple rotations. A recent analysis of data from 32 long-term experiments from North America and Europe [11] found that crop yields for maize and small grain cereals increased at higher values of an index used to measure rotational diversity. However, when fitting quadratic regression models of crop yield on their diversity measure and other controls, the coefficient corresponding to the squared diversity measure was negative for all 3 outcome crop classes (maize, small grain winter cereal and small grain spring cereal) suggesting diminishing returns with each additional unit of diversity.
@@
Due to experimental data limitations, recent metanalyses studying crop rotation diversification did not explore how the diversification benefit varied with each precrop and outcome crop combination [11, 5, 9, 39]. Our results suggest that the benefit of diversification varies considerably, and can even be negative (Table A8; Figure 5). For example, we estimate a negative effect of diversification in settings where soybean is the precrop and corn is the outcome crop whereas we find considerable diversification benefits when there is a legume precrop and spring wheat is the outcome crop (Table A8).
3.4 Study limitations
The causal inference approach used relies upon the assumption that there are no unmeasured confounders. While we controlled for potential confounders such as weather and geolocation, we did not control for soil quality in our analysis. It is possible that on more fertile land farmers are less likely to rotate their crops, because in such land high crop yields can be obtained without rotation. In Appendix A.1, we find that adding soil covariates from the USDA Soil Survey Geographic Database [40] as controls in the model barely changed our results. This sensitivity check was not conducted in Canada, China, and France due to lack of soil data, but our findings in the United States suggest that omission of soil covariates does not substantially bias our causal effect estimates and our estimates of how they vary with weather.
Nonetheless, our estimates for the benefits of rotation are still useful for understanding the effects of rotating crops on crop yield when using the typical reduction in fertilizer and pesticide use for a given rotation and region. For example, we still find a positive benefit of crop rotation even though less fertilizer and pesticide inputs are typically used in rotated fields. While our estimates do not capture the cost savings from using reduced chemical inputs, they can be coupled with work assessing the impact of rotation on chemical input use and how those impacts vary with weather to obtain a more complete picture of the comprehensive benefits of rotation and how they vary with weather. Furthermore, as shown mathematically in Appendix E, even when there are unmeasured mediators such as fertilizer and pesticide use, our estimates of how the rotation benefits vary with weather could (under certain assumptions) still give unbiased estimates of how the direct rotation benefits vary with weather.
Third, the outcome and treatment variables used do not perfectly capture the actual outcome and treatment variables of interest. The outcome variable is a remotely sensed greenness measure that is merely a proxy for crop yield. In Appendix F, we demonstrate that even though we do not use actual yield estimates as the outcome variable, our approach of using a vegetation index-based outcome variable coupled with an appropriate selection of weather covariates would give the same results, up to an unknown scaling factor, as the observational approach taken in recent works that assess the impacts of crop rotation [12] and tillage [43] on crop yield. Our rescaling of the estimated effects (Section 4.7) mitigates the discrepancy between our approach and those taken in [12, 43] by approximating the unknown scaling factor using subnational-level crop yield data. In Appendix F.5, we find that even when using modern, satellite-based yield estimates [44] instead of a vegetation index-based proxy as the outcome variable, the results still roughly differ by an unknown scaling factor, and moreover, for most rotations, county-level regressions of average estimated yield on the average of the proxy gave good approximations of the unknown scaling factors. In addition, in Appendix A.2, we find that our results were robust to the choice of vegetation index that was used as a yield proxy.
Fourth, our estimates for rotation benefits are not perfectly comparable across precrop choices. This is because even within the same country, we only estimated the effect of a rotation on the overlap region where both the treatment and control were commonly observed. For example, when comparing the SoybeanCorn rotation to the Winter WheatCorn rotation in the US, the estimates for the former are for parts of the US where both SoybeanCorn and CornCorn are commonly observed whereas the latter are for parts of the US where both Winter WheatCorn and CornCorn are commonly observed. A similar issue is that the rotation benefits cannot be directly compared across countries even for the same precrop and outcome crop choices because the analysis in different countries used crop type maps with different accuracies. For example, we expect rotation benefit estimates from France to suffer from the least attenuation bias compared to those in other countries due to the higher accuracy of crop maps in France (unlike in the US, Canada, and China the crop type maps in France were not based on remote sensing predictions).
4 Materials and Methods
4.1 Data sources and sample
@@
We used country-specific crop type maps (Table 1) to construct our study sample and to infer the crop type at each year and location within our study sample. These annual crop type maps had spatial resolutions of at most 30m 30m and gave the primary crop for each year and pixel, ignoring cover crops. Note that double cropping was not accounted for as a separate category in the crop type maps from China and Canada; however, this practice is considered uncommon in northeast China [51], Canada [52], and France [53]. We considered pixels that were classified as “double crop” in the US or as a “mixture” in France, as separate crop type categories, but these categories were not so common and thus were ultimately excluded from the analysis (Section 4.3).
@@
We used Google Earth Engine (GEE) and these country-specific crop type maps to construct a sample of geographical points corresponding to croplands. To include more samples with crop types of interest, we randomly sampled geographical points after masking for locations in which crop types of interest were observed in at least 20% of the years. Each country had a slightly different random sampling scheme due to differences across countries in the number of years with available crop type maps and in the crop types of interest (see Appendix B for details). For each country and sampled geographical point, the country-specific crop type maps were again used to infer the crop that was grown during every year with an available crop type classification.
@@
For the purposes of a sensitivity analysis (in Appendix A.2), at each location and year in our sample, we also extracted a time series of the near infrared reflectance of vegetation (NIRv) as an alternative to GCVI. NIRv, a vegetation index developed in [60], is the product of the Normalized Difference Vegetation Index and the NIR band, and NIRv has shown strong performance for predicting crop yields [61]. In particular, to extract NIRv values, we used the formula NIRv=NIR(NIRRED)/(NIRRED) where NIR and RED are bands from the multispectral satellite images.
The dataset was harmonized using latitude, longitude, and year. The satellite-based GCVI in each country had the same spatial resolution as the corresponding crop type map, and we extracted the GCVI data from the Landsat (or Sentinel-2) pixels that intersected our sampled pixels from the crop type maps. The raw weather dataset had a lower spatial resolution than the crop type maps, so the latitude and longitude of each pixel was used to determine the grid cell from which the weather data was extracted.
4.2 Normalized peak greenness as an outcome variable
@@
For some locations and years in our sample, we could not compute the peak GCVI value because the number of cloud free observations was too small to implement ordinary least squares regression. In other cases, likely also due to a dearth of cloud free observations, outlier peak GCVI values were observed, which we removed from the analysis (see Appendix C). Ultimately, when restricting our attention in each country to cases where the outcome crops of interest are grown and when averaging across each country, percent of observations were dropped for the former reason and an additional percent of samples were dropped for the latter reason.
@@
4.3 Types of rotations studied
@@
Samples where the precrop and outcome crop were the same were classified as having no rotation and were used as control units. For each country and outcome crop of interest, Figure 2 depicts the distribution of precrops in the sample. Samples where the outcome crop and precrop were different were classified as having a rotation; however, because we wanted to compare the effects of rotation across different precrop and outcome crop combinations, each precrop and outcome crop combination was categorized as a different type of rotation. In particular, a location and year in our sample with precrop A and outcome crop B was categorized as having an AB rotation. In order to focus our study on the most common crop rotations, we applied two selection criteria. For each country and outcome crop of interest within the country, we first only considered precrops that were among the 5 most prevalent precrops used in rotation. Second, we only considered precrops for which the number of samples with that precrop exceeded one tenth of the number of samples with the no rotation control. Table A8 shows the categories of rotation considered within each country after applying these selection criteria. It also shows the sample sizes of the treatment and control groups.
4.4 Causal forest analysis
@@
For each country, precrop A, and outcome crop B (corresponding to a unique row in Table A8), we conducted the following procedure to estimate the benefit of the AB rotation. First, we subsetted our data to only include samples with AB rotations or BB rotations and designated a binary treatment variable to be for samples with AB rotations and for samples with BB. Next we set the outcome variable to be the normalized peak GCVI variable described in Section 4.2. Finally we let denote a vector of year, latitude, longitude, 7 weather covariates, and (in the US only) irrigation status. The weather covariates captured early precipitation, growing season precipitation, average maximum and minimum daily temperatures in the growing season, and VPD for the three peak months of the growing season. In particular, when winter wheat was not the outcome crop (versus when winter wheat was the outcome crop) our 7 weather covariates were: total January-April (January-March) precipitation, total May-September (April-June) precipitation, average maximum daily temperature during April-September (April-June), average minimum daily temperature during April-September (April-June), June (April) VPD, July (May) VPD, and August (June) VPD. For samples in the US we also included irrigation status as an 11th covariate in , which we retrieved from the Landsat-based Irrigation Dataset [65] in the United States. Due to lack of estimates from this data product after the year 2017, and because irrigation status of fields do not typically change in the US, we imputed the irrigation status of each pixel after the year 2017 based on the proportion of years in which that pixel was classified as having irrigation between 2011 and 2017 (see Appendix D). Due to lack of irrigation data as well as greater homogeneity of irrigation status in the other study regions, we did not include irrigation as a covariate in Canada, France, or northeastern China.
@@
In words, if belongs to leaf , is the average of for samples with minus that for samples with when restricting our attention to the withheld samples belonging to leaf . The causal forest estimated the function via averaging honest trees according to . Note that when running the causal_forest function, we used the default setting of tune.parameters=“none”, meaning that no hyperparameters were tuned in the training of the causal forests and only prespecified or default hyperparameter values were used. Nonetheless, we found that the estimates of and produced were well calibrated and reasonably accurate (see Appendix I and Table A7).
@@
A standard way to summarize the treatment effects is the average treatment effect, given by ; however, estimation of this quantity is unstable in settings where some of the propensity scores are near or [66]. Because in our setting many samples had estimated propensity scores near or (Table A6), we instead estimated the overlap-weighted average treatment effect (ATO), given by
@@
The ATO is essentially a weighted average of the of the CATE that gives higher weight to points in regions of covariate space where the treatment and control are observed with similar frequency but no weight to points in regions of covariate space that have only treated units or only control units. The ATO was proposed in [67] and was shown in that paper to have the desirable property of minimizing asymptotic variance among a class of weighted average treatment effect estimands. We estimated the ATO using the average_treatment_effect function in the grf package. We briefly note that empirically, our ATO estimates were very similar to the mean estimated CATE when the mean is restricted to the sample of points with between 0.05 and 0.95 (Figure A13).
@@
We used the default settings of the average_treatment_effect function to calculate the standard errors of the estimated ATO. These standard errors accounted for the overlap-weighting scheme, but notably, as with all standard errors reported in the main text, they did not account for possible spatial correlations between the samples. In Appendix A.5, we justify our choice to not account for spatial correlations. In that appendix, we also show that the samples used in each analysis are relatively well spread out (median distance to nearest neighbor km) and conduct a sensitivity analysis where we find that accounting for spatial correlations increases the confidence interval widths by a factor of at most 1.7, but typically does so by a smaller factor (Figure A9).
@@
We computed the estimated ATO and its confidence interval using the above procedure for each rotation AB versus control BB comparison considered in Table A8. We interpreted the estimated ATO as an estimate for the average net effect of rotating using the precrop A on a proxy for the yield of crop B, where the net effect incorporates the direct effect of rotation and the indirect effect of rotation mediated by changes in fertilizer and pesticide input use. While is merely a vegetation index-based proxy for yield, the analysis in recent works such as [12, 43] which apply causal forests using yield estimates from the Scalable Crop Yield Mapper [59, 68, 69] would, up to a normalization constant, have given the same results as using a vegetation index under a principled choice of the vegatation index and weather controls (see Appendix F for a mathematical justification of this claim). Further, we convert estimates of the ATO to the crop yield scale in Section 4.7, which in essence approximates the unknown normalization constant using subnational-level crop yield data.
4.5 Analyzing heterogeneity of rotation benefits with weather

### v2

@@ 4.2 Discussion of heterogeneity with weather findings
While these findings of how rotation effects vary with weather are consistent with the existing literature (Appendix L), we note that our findings go far beyond the scope of most prior work. In particular, existing claims in the experimental literature looked at weather interactions after aggregating over many different precrop and outcome crop combinations, due to limitations in the number and types of experiments [10, 5, 9, 8]. Such claims can therefore reflect how precrops and outcome crops in experiments vary across regions with differing climates rather than how the benefit of rotation varies with weather. In addition, claims in the observational literature about how rotation benefits varied with weather all focused on a single country and only one of them provided uncertainty quantifications for these claims [12, 17, 19, 20] (Table S6).
@@ 4.3 Discussion of diversification benefit findings
Our results suggest that diversification beyond a simple rotation has diminishing returns (see Table S9 or Figures 3 and 5). In particular, we find that the average benefit of a 2-crop rotation across our 36 rotation-country pairs is about 3.7 times larger than the average additional benefit of switching from 2-crop rotations to a more diverse sequence. This finding agrees with a recent metanalysis of 45 experiments in China [9] that found crop rotation increased yields by 20 when compared to monoculture on average but extended (diversified) rotations provided only an additional 4 yield boost. Similarly, a recent global metanalysis of 462 crop rotation experiments involving legumes found that the inclusion of an additional legume precrop in a rotation had lower benefits when the control rotation was already diverse [5]. A recent analysis of field experiments in the European Union and Africa [56] found no additional benefits from increased diversification beyond a switch from monocropping to simple rotations. A recent analysis of data from 32 long-term experiments from North America and Europe [11] found that crop yields for maize and small grain cereals increased at higher values of an index used to measure rotational diversity. However, when fitting quadratic regression models of crop yield on their diversity measure and other controls, the coefficient corresponding to the squared diversity measure was negative for all 3 outcome crop classes (maize, small grain winter cereal and small grain spring cereal) suggesting diminishing returns with each additional unit of diversity.
@@
Due to experimental data limitations, recent metanalyses studying crop rotation diversification did not explore how the diversification benefit varied with each precrop and outcome crop combination [11, 5, 9, 56]. Our results suggest that the benefit of diversification varies considerably, and can even be negative (Table S9; Figure 5). For example, we estimate a negative effect of diversification in settings where soybean is the precrop and corn is the outcome crop whereas we find considerable diversification benefits when there is a legume precrop and spring wheat is the outcome crop (Table S9).
4.4 Limitations
This study has several limitations. First, our causal inference approach relies on the assumption that there are no unmeasured confounders. Due to data availability constraints we did not control for potential confounders such as soil quality (although, including soil variables in the US barley changed our results (Appendix C.1)). We also did not account for management practices such as tillage, cover cropping, residue management, fertilizer use, and pesticide use; however, these unmeasured management practices can be viewed as unmeasured mediators as opposed to unmeasured confounders, and we estimate the net rather than direct effects of crop sequence choice (see Section A.8). Moreover, farms that rotate tend to use less fertilizer [40] and pesticide [57], so our analysis does not capture the cost savings reductions in chemical inputs that are associated with rotations.
Second, we do not have exact measurements of the treatment and outcome variables which are instead estimated using satellite data. Misclassifications in the treatment variable likely lead to attenuation bias [12, 58], which as with the issue of unmeasured fertilizer and pesticide use, results in underestimation of rotation benefits. Appendix H suggests that our choice to use a vegetation index as the outcome variable rather than crop yield estimates from [30, 59] does not substantially influence the results and we also find robustness to the choice of vegetation index (Appendix C.2).
Third, the study has some limitations in its scope. We do not study the well-documented longer-term benefits of crop rotation [8, 11]. In addition, our analysis only considers the impact of a cropping sequence on the final crop in a cropping sequence rather than all crops in the sequence. Ultimately, farmers make decisions about which rotations to use based on many factors, including the expected yield and price of each of the crops in the rotation sequence; however, future work can couple our results and methods with crop rotation decision support systems [60, 61, 62, 63, 64].
A more comprehensive and detailed discussion of the limitations can be found in Appendix B. That appendix also summarizes sensitivity analyses and mathematical results, which assess whether or not the limitations discussed are severe.
5 Conclusion
Using satellite data and causal machine learning, we estimated the impacts of various cropping sequences and how they varied with weather. While randomized field experiments can reliably estimate the effects of cropping sequences on crop yield, given the limited geographical span of experimental data, drawing conclusions of how rotation benefits vary with weather remains a substantial challenge. Our results using observational approaches suggest that precrop effects are larger under rainier conditions. Moreover, the results suggest that if the temperature increases, precrop effects will decrease in cases where the precrop is a legume but increase in cases where the outcome crop is soybean. Estimates of how rotation benefits vary with weather can be coupled with climate models to forecast how the effectiveness of crop rotation could change in a changing climate (as is done for the Soybean-Corn rotation in [20]) and can inform agricultural adaptation to climate change.
In addition, our observational approach can uncover or fill in gaps in the experimental crop rotation literature. For example, as discussed in Section 4.1 our results for the SoybeanWinter Wheat sequence point to possible external validity issues regarding planting dates and weather conditions in the experimental literature. In addition, our approach can be used in any region with multiple years of crop type maps in order to identify cropping sequences that may be highly beneficial but have been understudied in randomized field experiments.
Acknowledgments
This work was supported by a Stanford Interdisciplinary Graduate Fellowship and the NASA Harvest and ACRES Consortia (NASA Applied Sciences Grant No. 80NSSC23M0032 (sub-award 125062-Z6521205) and 80NSSC23M0034 (sub-award 124245-Z6512205)). The authors wish to thank Kevin Guo and Stefan Wager for helpful discussions and anonymous reviewers for helpful comments.
Data and code availability
Supplement for “Precrop payoffs: causal machine learning reveals large but variable yield benefits of crop rotation in major breadbaskets”
Appendix A Materials and Methods
A.1 Data sources and sample
@@
We used country-specific crop type maps (Table S1) to construct our study sample and to infer the crop type at each year and location within our study sample. These annual crop type maps had spatial resolutions of at most 30m 30m and gave the primary crop for each year and pixel, ignoring cover crops. Note that double cropping was not accounted for as a separate category in the crop type maps from China and Canada; however, this practice is considered uncommon in northeast China [24], Canada [65], and France [66]. We considered pixels that were classified as “double crop” in the US or as a “mixture” in France, as separate crop type categories, but these categories were not so common and thus were ultimately excluded from the analysis (Section A.3).
@@
We used Google Earth Engine (GEE) and these country-specific crop type maps to construct a sample of geographical points corresponding to croplands. To include more samples with crop types of interest, we randomly sampled geographical points after masking for locations in which crop types of interest were observed in at least 20% of the years. Each country had a slightly different random sampling scheme due to differences across countries in the number of years with available crop type maps and in the crop types of interest (see Appendix D for details). For each country and sampled geographical point, the country-specific crop type maps were again used to infer the crop that was grown during every year with an available crop type classification.
For the purposes of a sensitivity analysis (in Appendix C.2), at each location and year in our sample, we also extracted a time series of the near infrared reflectance of vegetation (NIRv) as an alternative to GCVI. NIRv, a vegetation index developed in [67], is the product of the Normalized Difference Vegetation Index and the NIR band, and NIRv has shown strong performance for predicting crop yields [68]. In particular, to extract NIRv values, we used the formula NIRv=NIR(NIRRED)/(NIRRED) where NIR and RED are bands from the multispectral satellite images.
@@
The dataset was harmonized using latitude, longitude, and year. The raw TerraClimate weather dataset had a spatial resolution of ° ° (approximately 4km 4km) which was a substantially lower spatial resolution than that of the crop type maps. Therefore, the latitude and longitude of each sampled pixel from the crop type map was used to determine the grid cell from which the TerraClimate weather data was extracted. The satellite-based GCVI in each country had the same spatial resolution as the corresponding crop type map, and we extracted the GCVI data from the Landsat (or Sentinel-2) pixels that intersected our sampled pixels from the crop type maps.
A.2 Normalized peak greenness as an outcome variable
@@
For some locations and years in our sample, we could not compute the peak GCVI value because the number of cloud free observations was too small to implement ordinary least squares regression. In other cases, likely also due to a dearth of cloud free observations, outlier peak GCVI values were observed, which we removed from the analysis (see Appendix E). Ultimately, when restricting our attention in each country to cases where the outcome crops of interest are grown and when averaging across each country, percent of observations were dropped for the former reason and an additional percent of samples were dropped for the latter reason.
@@
A.3 Types of rotations studied
@@
Samples where the precrop and outcome crop were the same were classified as having no rotation and were used as control units. For each country and outcome crop of interest, Figure 2 depicts the distribution of precrops in the sample. Samples where the outcome crop and precrop were different were classified as having a rotation; however, because we wanted to compare the effects of rotation across different precrop and outcome crop combinations, each precrop and outcome crop combination was categorized as a different type of rotation. In particular, a location and year in our sample with precrop A and outcome crop B was categorized as having an AB rotation. In order to focus our study on the most common crop rotations, we applied two selection criteria. For each country and outcome crop of interest within the country, we first only considered precrops that were among the 5 most prevalent precrops used in rotation. Second, we only considered precrops for which the number of samples with that precrop exceeded one tenth of the number of samples with the no rotation control. Table S9 shows the categories of rotation considered within each country after applying these selection criteria. It also shows the sample sizes of the treatment and control groups.
A.4 Causal forest analysis
@@
For each country, precrop A, and outcome crop B (corresponding to a unique row in Table S9), we conducted the following procedure to estimate the benefit of the AB rotation. First, we subsetted our data to only include samples with AB rotations or BB rotations and designated a binary treatment variable to be for samples with AB rotations and for samples with BB. Next we set the outcome variable to be the normalized peak GCVI variable described in Section A.2. Finally we let denote a vector of year, latitude, longitude, 7 weather covariates, and (in the US only) irrigation status. The weather covariates captured early precipitation, growing season precipitation, average maximum and minimum daily temperatures in the growing season, and VPD for the three peak months of the growing season. In particular, when winter wheat was not the outcome crop (versus when winter wheat was the outcome crop) our 7 weather covariates were: total January-April (January-March) precipitation, total May-September (April-June) precipitation, average maximum daily temperature during April-September (April-June), average minimum daily temperature during April-September (April-June), June (April) VPD, July (May) VPD, and August (June) VPD. For samples in the US we also included irrigation status as an 11th covariate in , which we retrieved from the Landsat-based Irrigation Dataset [37] in the United States. Due to lack of estimates from this data product after the year 2017, and because irrigation status of fields do not typically change in the US, we imputed the irrigation status of each pixel after the year 2017 based on the proportion of years in which that pixel was classified as having irrigation between 2011 and 2017 (see Appendix F). Due to lack of irrigation data as well as greater homogeneity of irrigation status in the other study regions, we did not include irrigation as a covariate in Canada, France, or northeastern China.
@@
In words, if belongs to leaf , is the average of for samples with minus that for samples with when restricting our attention to the withheld samples belonging to leaf . The causal forest estimated the function via averaging honest trees according to . [22] provides theoretical guarantees that under certain technical conditions and if the sample size is large, the function returned by the causal forest gives a good estimate of the function defined in Equation (2). Note that when running the causal_forest function, we used the default setting of tune.parameters=“none”, meaning that no hyperparameters were tuned in the training of the causal forests and only prespecified or default hyperparameter values were used. Nonetheless, we found that the estimates of and produced were well calibrated and reasonably accurate (see Appendix K and Table S8).
@@
A standard way to summarize the treatment effects is the average treatment effect, given by ; however, estimation of this quantity is unstable in settings where some of the propensity scores are near or [69]. Because in our setting many samples had estimated propensity scores near or (Table S7), we instead estimated the overlap-weighted average treatment effect (ATO), given by
@@
The ATO is essentially a weighted average of the of the CATE that gives higher weight to points in regions of covariate space where the treatment and control are observed with similar frequency but no weight to points in regions of covariate space that have only treated units or only control units. The ATO was proposed in [70] and was shown in that paper to have the desirable property of minimizing asymptotic variance among a class of weighted average treatment effect estimands. We estimated the ATO using the average_treatment_effect function in the grf package. We briefly note that empirically, our ATO estimates were very similar to the mean estimated CATE when the mean is restricted to the sample of points with between 0.05 and 0.95 (Figure S15).
@@
We used the default settings of the average_treatment_effect function to calculate the standard errors of the estimated ATO. These standard errors accounted for the overlap-weighting scheme, but notably, as with all standard errors reported in the main text, they did not account for possible spatial correlations between the samples. In Appendix C.5, we justify our choice to not account for spatial correlations. In that appendix, we also show that the samples used in each analysis are relatively well spread out (median distance to nearest neighbor km) and conduct a sensitivity analysis where we find that accounting for spatial correlations increases the confidence interval widths by a factor of at most 1.7, but typically does so by a smaller factor (Figure S10).
@@
We computed the estimated ATO and its confidence interval using the above procedure for each rotation AB versus control BB comparison considered in Table S9. We interpreted the estimated ATO as an estimate for the average net effect of rotating using the precrop A on a proxy for the yield of crop B, where the net effect incorporates the direct effect of rotation and the indirect effect of rotation mediated by changes in fertilizer and pesticide input use. While is merely a vegetation index-based proxy for yield, the analysis in recent works such as [12, 35] which apply causal forests using yield estimates from the Scalable Crop Yield Mapper [30, 31, 71] would, up to a normalization constant, have given the same results as using a vegetation index under a principled choice of the vegatation index and weather controls (see Appendix H for a mathematical justification of this claim). Further, we convert estimates of the ATO to the crop yield scale in Section A.7, which in essence approximates the unknown normalization constant using subnational-level crop yield data.
A.5 Analyzing heterogeneity of rotation benefits with weather